OpenStudy (anonymous):

The value of y varies directly with x, and y = -3 when x = 11. Find y when x = 66.

jimthompson5910 (jim_thompson5910):

`The value of y varies directly with x` means y = k*x where k is some fixed number (a constant)

jimthompson5910 (jim_thompson5910):

we're told that ` y = -3 when x = 11` so we can use this to find k y = k*x -3 = k*11 ... plug in x = 11 and y = -3 solve for k